### v1

@@ 2 Methods
For each of the seven games, we obtain human-playing data from the MobLab Classroom economics experiment platform, which consists of 68,722 subjects from more than a hundred countries, spanning multiple years. The subjects are mostly, but not exclusively, college students who majored in social sciences. The individual responses for each game are recorded, creating a distribution of human behavior for that game. More details about the human-playing data are described in Sec. A.1 in the Supporting Information.
@@ 3.2 Deciphering Behavior: Prompts can be used to better understand individual human behaviors
$R^{2}$ = 0.43, MAE = 9.36 $R^{2}$ = 0.39, MAE = 5.86 $R^{2}$ = 0.45, MAE = 7.80
Intercept = 41.53 Intercept = 39.79 Intercept = 31.22
@@
generous 12.05 generous 5.58 pragmatic -5.47
generosity 9.28 generosity 4.52 rational -5.27
retain -8.98 gain -3.75 high 5.05
decision 6.27 create 3.39 standard 4.99
gain -6.13 highly 3.25 strong 3.73
focus -5.98 value 2.63 decision -3.69
offer -5.74 accept -2.63 significant 3.43
maximize -3.71 positive 2.37 advantageous 3.19
self -3.66 mutual 2.37 negotiation 3.18
goodwill 3.35 ensure 2.33 choice -3.10
@@
$R^{2}$ = 0.64, MAE = 11.10 $R^{2}$ = 0.51, MAE = 6.35 $R^{2}$ = 0.63, MAE = 1.82 $R^{2}$ = 0.67, MAE = 8.04
Intercept = 49.90 Intercept = 77.08 Intercept = 5.92 Intercept = 42.25
@@
risk -10.47 balance 8.05 benefit 2.77 risk -9.48
small -8.50 profit -4.40 resource -1.80 conservative -8.95
conservative -8.22 prioritize -3.64 collective 1.64 reward 7.19
maximize 8.01 fair 3.59 term 1.47 minimize -6.56
return 6.77 trust 3.59 group 1.32 maximize 5.78
substantial 6.53 maximize -3.41 long -1.22 prioritize -5.73
cautious -6.37 focus -3.12 strategy -1.21 safety -5.50
balance -6.18 self -2.85 potential -1.20 avoid -5.37
minimize -5.93 relationship 2.69 high 1.08 probability 4.81
significant 5.19 long 2.68 cooperation 1.06 cautious -4.79
Table 1: Linear regression analysis of elicited behaviors based on keywords in behavioral codes. Each behavioral code is converted into a 50-dimensional binary feature vector, representing keyword occurrences, to predict the mean behavior generated with that code.. This prediction is performed using an ordinary least square (OLS) linear regression model. 10 keywords with the highest absolute regression coefficients are listed for each game. The full regression table is provided in the Supporting Information (Table 5).
@@
To mitigate potential multicollinearity among keywords, we apply a Principal Component Analysis (PCA) to the 50-dimensional keyword vectors representing behavioral codes for each game. Our analysis reveals that the first few principal components exhibit a significant correlation with the behavioral choices in each game. As shown in Figure 3, for every game, at least one of the first three principal components of its behavioral codes has a moderate to strong correlation with the elicited behaviors ($|r|$ in between 0.438 to 0.618, $p<0.001$). The correlation of all first five principal components to the behavioral choices can be found in Figure 10 in the Supporting Information.
@@
We have seen how our deciphering process can be used to group and better understand games. Next, we use the codes to group and better understand distributions of human behaviors across games rather than just within them. Figure 5 demonstrates how the weighted codes found to generate distributions of human plays within games are laid out in the 2D projection map of all behavioral codes, when mixing across all games. The “activated” codes (with non-negligible weights $>0.001$) are not distributed evenly. In the four-game cluster, the activated codes have a high presence in areas according to behavioral strategies such as “assertive and self-advantaging decisions” and “highly competitive and profit-oriented”, a moderate presence in areas related to “fairness driven and collaborative”, “strategic generosity for high acceptance”, “rational and pragmatic”, and “balanced rationality and generosity”, and a low presence in areas related to “strategic balancing of fairness and profit”, “negotiation and relationship management”, and “generosity-oriented decision-making”. This characterization is consistent with perceptions about the behavioral tendencies of students (e.g., [5]), which are the major composite of our player population. In the Investor-Bomb games cluster, there is a concentration on two extremes, “conservative and risk-averse strategies” and “ambitious and high-risk, high-reward strategies”, and a low coverage in the middle ground, “moderate and risk-adjusted investing”, amid the high presence of “balanced and strategic decision-making” in the Bomb game. A similar pattern is observed in the banking cluster, where there lacks a middle ground strategy between “highly competitive and profit-driven banking” and “empathetic and fair-minded banking.” These behavioral markers provide a more informative and coherent characterization of the testing population than the distributions in Figure 9.
@@
Thus, behavioral codes can also serve as a tool for identifying distinct decision-making patterns across different populations. Figure 12 in the Supplementary Information displays the behavior signatures of various subject populations from the meta study of the Dictator Game by [6]. The five groupings include two labeled as students or non-students by [6], as well as three categorized by the type of country or ethnic group. We have labeled those three as ‘High Income,’ ‘Middle Income,’ and ‘Small Scale’. These names are changed from what [6] referred to as ‘Western’ (which include Germany, Sweden and the US, among others), ‘Developing’ (which include Russia, South Africa, and Honduras, among others) and ‘Primitive’ (which include the Tsimane, Hadza, and Mpakama, among others). Figure 6 highlights the weighted behavioral codes for these populations, revealing distinctive patterns in their decision-making tendencies: (i) In the ‘Student’ population, a substantial proportion of individuals exhibit behaviors concentrated in the ‘Highly Competitive and Profit-Oriented’ region. In contrast, the ‘Non-Student’ population demonstrates a stronger inclination toward ‘Generosity-Oriented Decision-Making’ as well as ‘Fairness-Driven and Collaborative’ choices. (ii) For subjects in High Income and Middle Income countries, ‘Highly Competitive and Profit-Oriented’ strategies are prominent, with the subjects from High Income societies showing a slight tendency toward ‘Generosity-Oriented Decision-Making’. By contrast, subjects from Small Scale populations exhibit a stronger emphasis on ‘Fairness-Driven and Collaborative’ behaviors, suggesting a greater inclination toward equitable resource distribution and cooperative decision-making. Our findings are aligned with the comparisons made between behaviors of these subject populations in the meta study [6] and provide interpretations of the motivations behind the revealed behavioral patterns.
@@ 4 Discussion
In our earlier work [7], we showed how games used to understand human behaviors could also be used as a Turing Test, to see if AI behaved similarly to humans, and to understand AI’s tendencies. Here, we have reversed that perspective. We have used AI to better understand human behavior, and also to better understand the various games and different types of settings in which humans interact. The strong fits and interpretability of the codes that we find suggest that this is a promising tool for understanding human behavior.
Our approach complements the range of approaches in the behavioral science literatures to understand and predict human behaviors, including various forms of revealed preferences where some form of objective or utility function is fit to best predict behaviors. This novel approach can be used as a general tool to facilitate behavioral science research in a variety of ways, such as creating virtual subjects and simulating experiments, screening potentially effective interventions; as well as designing, simulating and studying human-AI interactions. Future work can also utilize our approach to steer the behaviors of AI agents.
@@
Appendix A Methods
@@ A.1 Classic Economic Games and Human Behavior Data
For these seven game scenarios, we utilize human game-play data from Mei et al. [7] to conduct our experiments and analyses. This dataset was collected through the MobLab Classroom platform [^1] MobLab Classroom: https://www.moblab.com/products/classroom, retrieved 01/2025, covering a nine-year period from 2015 to 2023. The dataset comprises behavioral observations from 88,595 subjects across 15,236 sessions, showcasing significant geographical diversity with participants from 59 countries spanning multiple continents.
In our analyses, we focus exclusively on first-round play records within the games, yielding a refined subset of 68,772 subjects and 82,057 individual play records. This ensures that our analyses are based on initial, unconditioned decision-making behaviors.
@@ A.2 Deciphering Individual Behaviors
Game # of Learned Behavioral Codes
@@
Figure 7: Consistency of the behaviors elicited by the behavioral codes. For each deciphered behavioral code, we calculate the mean and standard deviation (std) of the behaviors that are elicited by the code. The mean and std are displayed correspondingly on the x- and y-axes.
@@ A.3 Distributional Alignment
Human Sample 1,000 random sample 0.70 0.70 0.68 1.05 0.56 1.05 0.87
LLM-default distribution Default system prompt 25.58 11.95 32.53 42.78 20.35 24.00 26.99
@@
The distributional alignment successfully fits the behavioral codes to a distribution of human play. Table 3 shows that the LLM-elicited distribution given the deciphered mixture of behavioral codes is indistinguishable from the human distribution in 5 out of 7 games, under a relaxed Kolmogorov-Smirnov (KS) test [^2] For the relaxed Kolmogorov-Smirnov (KS) test, samples are binned with a width of 10 (the bin width for the Public Goods game is 2)..
@@ Appendix D Population Signatures
For the MobLab player population, the weighted behavioral codes are highlighted on the 2D projection map shown in Figure 5. Semantic clusters on the map are summarized into keywords using ChatGPT. The projection reveals that the weighted codes are unevenly distributed across the space, offering detailed and nuanced insights into the characteristics of the player population.
Behavioral codes can also serve as a valuable tool for identifying distinct decision-making patterns across different populations. In a meta-study of the Dictator Game [6] [^3] Data from Engel [6]: https://osf.io/xc73h/, retrieved on March 14, 2025., various populations were analyzed, revealing distinct behavioral distributions. Figure 12 displays the behavior distributions of various populations in the Dictator game, including comparisons between Students and Non-students populations, as well as populations from diverse societal backgrounds. Figure 6 highlights the weighted behavioral codes for these populations, revealing distinctive patterns in their decision-making tendencies: (i) In the Student population, a substantial proportion of individuals exhibit behaviors concentrated in the ‘Highly Competitive and Profit-Oriented’ region. In contrast, the Non-Student population demonstrates a stronger inclination toward ‘Generosity-Oriented Decision-Making’ as well as ‘Fairness-Driven and Collaborative’ choices. (ii) In both the High and Middle Income populations, ‘Highly Competitive and Profit-Oriented’ strategies are prominent, with the High Income population showing a slight tendency toward ‘Generosity-Oriented Decision-Making’. By contrast, the Small-Scale population exhibits a stronger emphasis on ‘Fairness-Driven and Collaborative’ behaviors, suggesting a greater inclination toward equitable resource distribution and cooperative decision-making.

### v2

@@ 2 Methods
For each of the seven games, we obtain human-playing data from the MobLab Classroom economics experiment platform, which consists of 68,779 subjects from more than a hundred countries, spanning multiple years. The subjects are mostly, but not exclusively, college students who majored in social sciences. The individual responses for each game are recorded, creating a distribution of human behavior for that game. More details about the human-playing data are described in Sec. A.1 in the Supporting Information.
@@ 3.2 Deciphering Behavior: Prompts can be used to better understand individual human behaviors
$R^{2}$ = 0.43, MAE = 9.35 $R^{2}$ = 0.39, MAE = 5.86 $R^{2}$ = 0.45, MAE = 7.80
Intercept = 41.55 Intercept = 39.78 Intercept = 31.17
@@
generous 12.05 generous 5.57 pragmatic -5.49
generosity 9.28 generosity 4.53 rational -5.28
retain -8.95 gain -3.74 high 5.03
decision 6.26 create 3.40 standard 5.01
gain -6.13 highly 3.25 strong 3.74
focus -6.00 accept -2.64 decision -3.60
offer -5.75 value 2.62 significant 3.42
maximize -3.71 mutual 2.37 advantageous 3.18
self -3.65 positive 2.37 negotiation 3.17
goodwill 3.37 ensure 2.33 choice -3.09
@@
$R^{2}$ = 0.64, MAE = 11.05 $R^{2}$ = 0.51, MAE = 6.35 $R^{2}$ = 0.63, MAE = 1.82 $R^{2}$ = 0.67, MAE = 8.01
Intercept = 50.19 Intercept = 76.96 Intercept = 5.93 Intercept = 42.27
@@
risk -10.71 balance 8.07 benefit 2.76 risk -9.51
small -8.68 profit -4.42 resource -1.80 conservative -8.93
conservative -8.17 prioritize -3.64 collective 1.64 reward 7.24
maximize 8.02 fair 3.61 term 1.48 minimize -6.58
return 6.67 trust 3.61 group 1.32 maximize 5.85
substantial 6.53 maximize -3.39 long -1.22 prioritize -5.69
cautious -6.42 focus -3.15 potential -1.20 safety -5.42
balance -6.06 self -2.83 strategy -1.20 avoid -5.40
minimize -5.91 relationship 2.72 high 1.08 cautious -4.79
significant 5.19 long 2.65 cooperation 1.06 probability 4.75
Table 1: Linear regression analysis of elicited behaviors based on keywords in behavioral codes. Each behavioral code is converted into a 50-dimensional binary feature vector, representing keyword occurrences, to predict the mean behavior generated with that code. This prediction is performed using an ordinary least square (OLS) linear regression model. 10 keywords with the highest absolute regression coefficients are listed for each game. The full regression table is provided in the Supporting Information (Table 5).
@@
To mitigate potential multicollinearity among keywords, we apply a Principal Component Analysis (PCA) to the 50-dimensional keyword vectors representing behavioral codes for each game. Our analysis reveals that the first few principal components exhibit a significant correlation with the behavioral choices in each game. As shown in Figure 3, for every game, at least one of the first two principal components of its behavioral codes has a moderate to strong correlation with the elicited behaviors ($|r|$ in between 0.432 to 0.620, $p<0.001$). The correlation of all first five principal components to the behavioral choices can be found in Figure 10 in the Supporting Information.
@@
We have seen how our deciphering process can be used to group and better understand games. Next, we use the codes to group and better understand distributions of human behaviors across games rather than just within them. Figure 5 demonstrates how the weighted codes found to generate distributions of human plays within games are laid out in the 2D projection map of all behavioral codes, when mixing across all games. The “activated” codes (with non-negligible weights $>0.001$) are not distributed evenly. In the four-game cluster, the activated codes have a high presence in areas according to behavioral strategies such as “Selfish Maximization Tactics” and “High-Value Negotiation”, a moderate presence in areas related to “Diplomatic Fairness Strategy”, “Generous Negotiation Strategy”, “Rational Acceptance Threshold”, and “Balanced Negotiation Offers”, and a low presence in areas related to “Balanced Cooperative Gains”, “Fair Profit Balance”, and “Generous Resource Sharing”. This characterization is consistent with perceptions about the behavioral tendencies of students (e.g., [5]), which are the major composite of our player population. In the Investor-Bomb games cluster, there is a concentration on two ends, “Risk-Averse Investing” and “Assertive Cautious Investing”, and a low coverage in the middle ground, “Moderate Investment Strategy”, amid the high presence of “Risk-Reward Balancing” in the Bomb game. A similar pattern is observed in the banking cluster, where there lacks a middle ground strategy between “Profit-Maximizing Banker” and “Cooperative Banker Tactics.” These behavioral markers provide a more informative and coherent characterization of the testing population than the distributions in Figure 9. The generation process of the cluster labels is detailed in Appendix D.
@@
Thus, behavioral codes can also serve as a tool for identifying distinct decision-making patterns across different populations. Figure 12 in the Supplementary Information displays the behavior signatures of various subject populations from the meta study of the Dictator Game by [6]. The five groupings include two labeled as students or non-students by [6], as well as three categorized by the type of country or ethnic group. We have labeled those three as ‘High Income,’ ‘Middle Income,’ and ‘Small Scale’. These names are changed from what [6] referred to as ‘Western’ (which include Germany, Sweden and the US, among others), ‘Developing’ (which include Russia, South Africa, and Honduras, among others) and ‘Primitive’ (which include the Tsimane, Hadza, and Mpakama, among others). Figure 6 highlights the weighted behavioral codes for these populations, revealing distinctive patterns in their decision-making tendencies: (i) In the ‘Student’ population, a substantial proportion of individuals exhibit behaviors concentrated in the “Selfish Maximization Tactics” region. In contrast, the ‘Non-Student’ population demonstrates a stronger inclination toward “Generous Resource Sharing” as well as “Diplomatic Fairness Strategy” choices. (ii) For subjects in High Income and Middle Income countries, “Selfish Maximization Tactics” strategies are prominent, with the subjects from High Income societies showing a slight tendency toward “Generous Resource Sharing”. By contrast, subjects from Small Scale populations exhibit a stronger emphasis on “Diplomatic Fairness Strategy” behaviors, suggesting a greater inclination toward equitable resource distribution and cooperative decision-making. Our findings are aligned with the comparisons made between behaviors of these subject populations in the meta study [6] and provide interpretations of the motivations behind the revealed behavioral patterns.
@@ 4 Discussion
In our earlier work [7], we showed how games used to understand human behaviors could also be used as a Turing Test, to see if AI behaved similarly to humans, and to understand AI’s tendencies. Here, we have reversed that perspective. We have used AI to better understand the motivations behind human behavior, and also to better understand the various games and different types of settings in which humans interact. The strong fits and interpretability of the codes that we find suggest that this is a promising tool for understanding human behavior.
Our approach complements existing approaches in the behavioral sciences used to understand and predict human behaviors, including various forms of revealed preferences where some form of objective or utility function is fit to best predict behaviors. Our approach features interpretability, extendability, and scalability; and it can be used as a general tool to facilitate behavioral science research in a variety of ways, such as creating virtual subjects and simulating experiments, screening potentially effective interventions; as well as designing, simulating and studying human-AI interactions. Future work can build upon our approach to steer the behaviors of AI agents and further investigate the cognitive processes underlying human decision-making.
@@
Data and Software Availability. Processed data files, code, and prompts will be deposited on Github upon publication.
Appendix A Supplementary Information
@@ A.1 Classic Economic Games and Human Behavior Data
For these seven game scenarios, we utilize human game-play data from Mei et al. [7] to conduct our experiments and analyses. This dataset was collected through the MobLab Classroom platform [^1] MobLab Classroom: https://www.moblab.com/products/classroom, retrieved 01/2025, covering a nine-year period from 2015 to 2023. The dataset comprises behavioral observations from 88,595 subjects across 15,236 sessions, showcasing significant geographical diversity with participants from 59 countries spanning multiple continents. In all the games described in this paper, student input is collected through an interactive slider bar on the MobLab platform, which allows participants to select a value along a predefined range. Once the desired value is set, students submit their choice.
In our analyses, we focus exclusively on first-round play records within the games, yielding a refined subset of 68,779 subjects and 82,057 individual play records. This ensures that our analyses are based on initial, unconditioned decision-making behaviors.
@@ A.2 Deciphering Individual Behaviors
Game # of Behavioral Codes
@@
Figure 7: Consistency of the behaviors elicited by the behavioral codes. For each deciphered behavioral code, we calculate the mean and standard deviation (std) of the behaviors that are elicited by the code. The mean and std are displayed correspondingly on the x- and y-axes. Note that the y-axis (std) display only 20% of the behavior range.
@@ A.3 Distributional Alignment
Human Sample 1,000 random sample 0.69 0.72 0.70 1.11 0.85 0.21 0.90
LLM-default distribution Default system prompt 25.58 11.35 33.99 28.23 20.00 25.49 16.62
@@
The distributional alignment successfully fits the behavioral codes to a distribution of human play. Table 3 shows that the LLM-elicited distribution given the deciphered mixture of behavioral codes is indistinguishable from the human distribution in 5 out of 7 games, under a relaxed Kolmogorov-Smirnov (KS) test [^2] In the relaxed Kolmogorov-Smirnov (KS) test, samples are binned with a width of 10 (the bin width for the Public Goods game is 2)..
@@ Appendix D Population Signatures
For the MobLab player population, the weighted behavioral codes are highlighted on the 2D projection map shown in Figure 5. Semantic clusters on the map are summarized into keywords using ChatGPT with the prompt below. The projection reveals that the weighted codes are unevenly distributed across the space, offering detailed and nuanced insights into the characteristics of the player population.
Here are 23 clusters of behavioral codes in the provided file. A behavioral code is a piece of text that induces behaviors with a natural language description of motivation or context. Please summarize these behavioral code clusters in to short titles (e.g., within four words).
The summarized cluster titles should be: (1) understandable, (2) semantically relevant to the behavioral codes inside the clusters, (3) well covering the whole cluster, (4) distinguishable from other clusters, (5) generate a generalizable cluster title and avoid including any information specific to particular game instructions (will be given below).
Before output, please go back and forth review and refine the summarized cluster titles so that they meet the requirements. Please output the cluster names in lines, followed by the summarized titles (e.g., “cluster 8AE92B-496: …”).
Game instructions: {game_instructions}
Behavioral codes can also serve as a valuable tool for identifying distinct decision-making patterns across different populations. In a meta-study of the Dictator Game [6] [^3] Data from Engel [6]: https://osf.io/xc73h/, retrieved on March 14, 2025., various populations were analyzed, revealing distinct behavioral distributions. Figure 12 displays the behavior distributions of various populations in the Dictator game, including comparisons between Students and Non-students populations, as well as populations from diverse societal backgrounds. Figure 6 highlights the weighted behavioral codes for these populations, revealing distinctive patterns in their decision-making tendencies.